OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair of points. (-15,-14) and (-19,-20)

hero (hero):

just insert each point into y = mx + b: -14 = -15m + b -20 =-19m + b Isolate b: b = 15m - 14 b = 19m - 20 Set b = b 15m - 14 = 19m - 20 Put like terms on the same side: 19m - 15m = 14 - 20 4m = -6 m = -6/4 m = -3/2

OpenStudy (anonymous):

Alright, thanks.